# On classes of bounded tree rank, their interpretations, and efficient sparsification

### v1

@@ Abstract
{textblock}
20(-1.7, 8.6) [Uncaptioned image] {textblock}20(-1.9, 8.9) [Uncaptioned image]
@@ 2 Preliminaries
Graph theory. We use $[k]$ to denote the set $\{1,\ldots,k\}$. We use mostly standard graph theoretic notation. Let $G$ be a graph. We write $N_{r}^{G}(v)$ for the closed $r$-neighborhood of a vertex $v$, that is, $N_{r}^{G}(v)$ is set of all vertices that are reachable from $v$ by a path with at most $r$ edges, including $v$. The distance between vertices $u$ and $v$, denoted by $dist_{G}(u,v)$, is the minimum number of edges in a path between $u$ and $v$. The radius of a graph is smallest integer $r$ so that there exists a vertex that has distance at most $r$ from every other vertex. Given a graph $G$ and a set $X\subseteq V(G)$, we write $G[X]$ for the subgraph of $G$ induced on $X$, and $G-X$ for the subgraph of $G$ induced on $V(G)\setminus X$.
@@ 3 Vertex rankings in sparse graphs
In this section we introduce a way to assign, for given parameters $r$ and $m$, to every vertex of a graph its $(r,m)$-rank. Intuitively, the $(r,m)$-rank of a vertex measures how complicated the $r$-neighborhood of $v$ is. Our definition is algorithmic and is given in Section 3.1 together with the proof that the $(r,m)$–rank can be computed in FPT runtime with respect to the parameters $r$ and $m$. Then we prove Theorems 6 and 7 relating graph classes for which suitable rankings of vertices exist to graph classes of bounded tree rank (Section 3.2) and bounded expansion (Section 3.3).
@@ 3.1 The ranking algorithm
Let $G$ be a graph and $r,m$ be positive integers. The $(r,m)$-ranking algorithm works as follows. Initially, each vertex is assigned rank $\infty$. Then the algorithm proceeds in rounds for $i=1,2,3,\ldots$ as follows. In the $i$-th round, the algorithm considers all vertices of rank $\infty$ (these are the vertices that have not received a finite rank in rounds $11,\ldots,i-$ 1), and for each such vertex $v$ it checks (in parallel) the following condition: Does there exist a set $S\subseteq V(G)\setminus\{v\}$ of size at most $m$ such that $N_{r}^{G-S}(v)\setminus\{v\}$ contains only vertices of finite rank? If yes, then $v$ receives rank $i$, otherwise it keeps rank $\infty$. (Notice that in the first round the algorithm assigns rank 1 to vertices of degree at most $m$.) The algorithm stops when all vertices obtain a finite rank or when in some round no new vertex receives a rank. Thus, in any case, the algorithm stops after at most $|V(G)|$ rounds.
@@ 3.1.1 Algorithmic considerations
Notice that the straightforward implementation of the ranking algorithm runs in time $\mathcal{O}(n^{m+4})$: we have at most $n$ rounds, in each round we consider at most $n$ vertices, and for each vertex we can in time $\mathcal{O}(n^{m+2})$ try removing all sets $S$ of size $m$ in time $\mathcal{O}(n^{m})$ and check whether $G-S$ satisfies the desired condition in time $\mathcal{O}(n^{2})$. Thus, we trivially get an XP algorithm. However, we can replace the subroutine which checks whether there exists a set $S$ with $|S|\leq m$ with desired properties by a simple branch-and-bound procedure with FPT runtime.
@@
The algorithm checks whether there is a path of length at most $r$ from $v$ to any vertex in $A$. If no such path exists, the algorithm outputs YES. If such a path $P$ exists and $m=0$, the algorithm outputs NO. If $m>0$ then at least one of the vertices on the path $P$ has to be in the solution $S$, and so we can branch on all vertices $u\in V(P)\setminus\{v\}$ (there are at most $r$ of them) and call the algorithm with input $G-\{u\},v,A\setminus\{u\}$, $r$, $m-1$. Then the algorithm returns YES if for at least one such $u$ the recursive call returned YES, and returns NO otherwise.
@@ 3.2 Vertex rankings and bounded tree rank
Let $T$ be a subgraph of $G$ that corresponds to an $\leq r$-subdivision of $T_{d,m+1}$. We prove that the principal vertices of $T$ of height $i$ (where the height is measured in $T_{d,m+1}$, and we think of leaves as having height 00) have $(r,m)$-rank at least $i+1$ in $G$. For $i=0$, the leaves of $T$ clearly have rank at least $1$. For $i>1$, let $v$ be a vertex of $T$ corresponding to a vertex of height $i$ in $T_{d,m+1}$. Then there are $m+1$ internally disjoint paths connecting $v$ to the corresponding $m+1$ children that have rank at least $i$. Therefore we cannot disconnect $v$ from all of these children by deleting $m$ vertices other than $v$. Thus $v$ has $(r,m)$-rank at least $i+1$, which completes the proof. ∎
@@
For $d=1$ we can set $m^{\prime}=m-1$. Then if there exists a vertex of $(r,m^{\prime})$-rank more than $1$ in $G$, this vertex has at least $m$ neighbors, and thus is the root of a $T_{1,m}$ subgraph in $G$. For $d>1$ we proceed as follows. For a given $m$, we wish to define a suitable $m^{\prime}$. First, let $W_{d,m,r}$ denote the number of vertices of the graph obtained from $T_{d-1,m}$ by subdividing each edge $r$ times. For convenience, set $M=m\cdot W_{d,m,r}+r\cdot m+m$. Let $m^{\prime\prime}$ be the number obtained by the inductive assumption applied to $d-1$, $r$, and $M$. Then we define $m^{\prime}:=\max\{m^{\prime\prime},r\cdot m\}$. Let $v$ be a vertex of $(r,m^{\prime})$-rank more than $d$ in $G$. Then for every set $S\subseteq V(G)\setminus\{v\}$ of size at most $m^{\prime}$, the set $N_{r}^{G-S}(v)\setminus\{v\}$ contains at least one vertex of $(r,m^{\prime})$-rank at least $d$.
We claim that there exist paths $P_{1},P_{2},\ldots,P_{m}$ of length at most $r$ that are disjoint other than at $v$, and so that each path $P_{i}$ joins $v$ to a vertex $u_{i}\neq v$ that has $(r,m^{\prime})$-rank at least $d$. We find the paths greedily by adding one path at a time. Suppose that so far we have found the paths $P_{1},\ldots,P_{j}$ for some $j<{m}$. Let $S$ be the set of all vertices $u\neq v$ that are in any of the paths $P_{1},\ldots,P_{j}$. This set $S$ contains at most $r$ vertices from each of the paths, and thus $S$ has size at most $r(m-1)\leq{m^{\prime}}$. Thus the set $N_{r}^{G-S}(v)\setminus\{v\}$ contains at least one vertex of $(r,m^{\prime})$-rank at least $d$, and we can add another path $P_{j+1}$ to our collection. This proves the claim. We note that this argument amounts to a Menger-type result about short paths, and this type of result has previously appeared in [16].
Since $m^{\prime}\geq m^{\prime\prime}$, the $(r,m^{\prime\prime})$-rank of each vertex of $G$ is at least its $(r,m^{\prime})$-rank. So by the inductive assumption applied to $d-1$, $r$, and $M$, each vertex $u_{i}$ is the root of an $\leq r$-subdivision of $T_{d-1,M}$. Let us call this $\leq r$-subdivision rooted at $u_{i}$ by $T_{i}$.
@@ 3.3 Vertex rankings and bounded expansion
Let $G$ be a graph with $\mathrm{scol}_{r}(G)=m$. Then $G$ has finite $(r,m-1)$-rank.
We stress that in Lemma 20 we are not bounding the maximum value of $(r,m-1)$-rank of any vertex of $G$, but merely claiming that the $(r,m-1)$-ranking algorithm will assign to each vertex of $G$ a finite value. This value, however, can be arbitrarily large (and in particular it is not bounded in terms of $r$ and $m$). This can be seen already on the class of trees – the class of all trees has bounded strong coloring numbers (one can set $m:=r+1$), but on the complete $m$-ary tree of depth $d$ the $(r,m-1)$-ranking algorithm will need $d$ rounds to finish.
@@
Assume for contradiction that there exists a vertex of $G$ of $(r,m-1)$-rank $\infty$. Fix an order $\leq$ on $V(G)$ that certifies that $\mathrm{scol}_{r}(G)=m$. Let $v$ be the smallest vertex of $G$ with respect to $\leq$ that has $(r,m-1)$-rank $\infty$. Let $A$ be the set of all vertices below $v$ in $\leq$, and let $i$ be the maximum $(r,m-1)$-rank of any vertex in $A$. We claim that in round $(i+1)$ of the ranking algorithm vertex $v$ received rank (meaning it has rank $i+1$), which is a contradiction with our assumption on $v$. To see this, let $S$ be the set of all vertices that are strongly $r$-reachable from $v$ in $G$. Note that every path of length at most $r$ from $v$ to a vertex $w\geq v$ must contain a vertex in $S$. It follows that $N_{r}^{G-S}(v)\setminus\{v\}$ contains only vertices of finite rank, a contradiction. ∎
@@ 4 A lemma about edge-stable graphs
Set $X_{t}:=S$. To show the existence of the remaining vertices $w_{1},\ldots,w_{t-1}$ and sets $X_{1},\ldots,X_{t-1}$, we will use the induction hypothesis with $t-1$ applied to the path $P^{\prime}:=v_{1}\ldots v_{q}$ and the set $S^{\prime}:=S\cap N^{G}(v_{q})$ (this is justified since $S^{\prime}\subseteq N^{G}(v_{q})$, the size of $S^{\prime}$ is as required, and $S^{\prime}\subseteq S$, so $S^{\prime}$ is disjoint from $N^{G}(v_{1})$). This yields vertices $w_{1},\ldots,w_{t-1}$ and sets $X_{1},\ldots,X_{t-1}$ with properties 1-3 above satisfied for $P^{\prime}$ and $S^{\prime}$. We now verify that $w_{1},\ldots,w_{t}$ together with $X_{1},\ldots,X_{t}$ satisfy properties 1-3 for $P$ and $S$. Let $i\in[t]$ be arbitrary.
@@ 5.1 Classes of tree rank at most 2
For the other direction, suppose that for every $r\in\mathbb{N}$ there exists $m\in\mathbb{N}$ such that every graph in $\mathcal{C}$ has $(r,m)$-rank at most $2$. Let $r\in\mathbb{N}$, and set $f(r):=(m+1)^{r}$ and $d(r):=m$. Let $v\in V(G)$ be arbitrary, and let $S$ be the set of vertices with $(r,m)$-rank exactly $2$ in $N_{r}^{G}(v)$ (these are the vertices of degree more than $m$). Going for a contradiction, we may assume that $|S|>f(r)$, since otherwise we are done. Thus, by Lemma 30 applied to $r$, $t:=m+1$ and $S$ in the subgraph $H$ of $G$ induced on $N_{r}^{G}(v)$, there exists a vertex $u\in N_{r}^{G}(v)$ such that there are $m+1$ vertices in $S$ that are distinct from $u$ and can be reached from $u$ by internally vertex-disjoint paths of length at most $r$. As all of the vertices in $S$ have $(r,m)$-rank $2$, this shows that the $(r,m)$-rank of $u$ is more than $2$, which is a contradiction. Thus $\mathcal{C}$ is of locally almost bounded degree with functions $f$ and $d$, as desired. ∎
@@ 5.2 The case of interpretations of bounded range
Assume for contradiction that such a graph $G\in\mathcal{C}$ exists. Let $u_{1},\ldots,u_{t},w_{1},\ldots,w_{t}$ be the vertices of a half-graph of order $t$ that $G$ contains as a semi-induced subgraph. Then all of these vertices $u_{1},\ldots,u_{t},w_{1},\ldots,w_{t}$ are in $N_{2}^{G}(u_{1})$. Also, for any $i,j\in\{1,2,\ldots,t\}$ with $i-j>k(2)$, we know that $u_{i}$ and $u_{j}$ are not $k(2)$-near twins, since their adjacency differs in vertices $w_{j},w_{j+1},\ldots,w_{i-1}$. Therefore, the vertices $u_{1},u_{k(2)+2},u_{2k(2)+3},\ldots u_{m(2)\cdot k(2)+m(2)+1}$ form a set of $m(2)+1$ pairwise not $k(2)$-near-twins, a contradiction with $\mathcal{C}$ being locally almost near-covered. ∎
@@
    every vertex of $A$ is adjacent to all but at most $2k$ vertices of $B$ and every vertex of $B$ is adjacent to all but at most $2k$ vertices of $A$
@@
[…]
@@
Now let us define an auxiliary graph $\widehat{G}$ where each vertex of $\widehat{G}$ is a tuple of the form $(A_{i},B_{i})$, and two distinct vertices $(A_{i},B_{i})$ and $(A_{j},B_{j})$ are adjacent if $A_{i}\cup B_{i}$ has non-empty intersection with $A_{j}\cup B_{j}$. So this graph $\widehat{G}$ has exactly $t$ vertices. Moreover, it has maximum degree less than $4(m(3)+1)$; for each vertex $(A_{i},B_{i})$, there are at most $12m(3)+$ 1 ways of choosing a tuple $(A,B)$ of mutually heavy parts of $\mathcal{F}$ so that either $A=A_{i}$ or $B=A_{i}$, and likewise with $B_{i}$ in place of $A_{i}$. This is using Claim Claim which implies that each vertex of $G/\mathcal{F}$ is mutually heavy with at most $m(3)$ other parts of $\mathcal{F}$ (and the extra $+1$ comes from the fact that a vertex may be mutually heavy with itself).
@@
[…]
@@
[…]

### v2

@@ 2 Preliminaries
Graph theory. We use $[k]$ to denote the set $\{1,\ldots,k\}$. We use mostly standard graph theoretic notation. Let $G$ be a graph. We write $N_{r}^{G}(v)$ for the closed $r$-neighborhood of a vertex $v$, that is, $N_{r}^{G}(v)$ is set of all vertices that are reachable from $v$ by a path with at most $r$ edges, including $v$. The distance between vertices $u$ and $v$, denoted by $dist_{G}(u,v)$, is the minimum number of edges in a path between $u$ and $v$. The radius of a graph is smallest integer $r$ so that there exists a vertex that has distance at most $r$ from every other vertex. Given a graph $G$ and a set $X\subseteq V(G)$, we write $G[X]$ for the subgraph of $G$ induced on $X$, and GXG-X for the subgraph of $G$ induced on $V(G)\setminus X$.
@@ 3 Vertex rankings in sparse graphs
In this section we introduce a way to assign, for given parameters $r$ and $m$, to every vertex of a graph its $(r,m)$-rank. Intuitively, the $(r,m)$-rank of a vertex measures how complicated the $r$-neighborhood of $v$ is. Our definition is algorithmic and is given in Section 3.1 together with the proof that the (r,m)(r,m)–rank can be computed in FPT runtime with respect to the parameters $r$ and $m$. Then we prove Theorems 6 and 7 relating graph classes for which suitable rankings of vertices exist to graph classes of bounded tree rank (Section 3.2) and bounded expansion (Section 3.3).
@@ 3.1 The ranking algorithm
Let $G$ be a graph and $r,m$ be positive integers. The $(r,m)$-ranking algorithm works as follows. Initially, each vertex is assigned rank $\infty$. Then the algorithm proceeds in rounds for $i=1,2,3,\ldots$ as follows. In the $i$-th round, the algorithm considers all vertices of rank $\infty$ (these are the vertices that have not received a finite rank in rounds $1,\ldots,i-1$), and for each such vertex $v$ it checks (in parallel) the following condition: Does there exist a set $S\subseteq V(G)\setminus\{v\}$ of size at most $m$ such that $N_{r}^{G-S}(v)\setminus\{v\}$ contains only vertices of finite rank? If yes, then $v$ receives rank $i$, otherwise it keeps rank $\infty$. (Notice that in the first round the algorithm assigns rank 1 to vertices of degree at most $m$.) The algorithm stops when all vertices obtain a finite rank or when in some round no new vertex receives a rank. Thus, in any case, the algorithm stops after at most $|V(G)|$ rounds.
@@ 3.1.1 Algorithmic considerations
Notice that the straightforward implementation of the ranking algorithm runs in time $\mathcal{O}(n^{m+4})$: we have at most $n$ rounds, in each round we consider at most $n$ vertices, and for each vertex we can in time $\mathcal{O}(n^{m+2})$ try removing all sets $S$ of size $m$ in time $\mathcal{O}(n^{m})$ and check whether GSG-S satisfies the desired condition in time $\mathcal{O}(n^{2})$. Thus, we trivially get an XP algorithm. However, we can replace the subroutine which checks whether there exists a set $S$ with $|S|\leq m$ with desired properties by a simple branch-and-bound procedure with FPT runtime.
@@
The algorithm checks whether there is a path of length at most $r$ from $v$ to any vertex in $A$. If no such path exists, the algorithm outputs YES. If such a path $P$ exists and $m=0$, the algorithm outputs NO. If $m>0$ then at least one of the vertices on the path $P$ has to be in the solution $S$, and so we can branch on all vertices $u\in V(P)\setminus\{v\}$ (there are at most $r$ of them) and call the algorithm with input $G-\{u\},v,A\setminus\{u\}$, $r$, m1m-1. Then the algorithm returns YES if for at least one such $u$ the recursive call returned YES, and returns NO otherwise.
@@ 3.2 Vertex rankings and bounded tree rank
Let $T$ be a subgraph of $G$ that corresponds to an $\leq r$-subdivision of $T_{d,m+1}$. We prove that the principal vertices of $T$ of height $i$ (where the height is measured in $T_{d,m+1}$, and we think of leaves as having height 0) have $(r,m)$-rank at least $i+1$ in $G$. For $i=0$, the leaves of $T$ clearly have rank at least $1$. For $i>1$, let $v$ be a vertex of $T$ corresponding to a vertex of height $i$ in $T_{d,m+1}$. Then there are $m+1$ internally disjoint paths connecting $v$ to the corresponding $m+1$ children that have rank at least $i$. Therefore we cannot disconnect $v$ from all of these children by deleting $m$ vertices other than $v$. Thus $v$ has $(r,m)$-rank at least $i+1$, which completes the proof. ∎
@@
For $d=1$ we can set $m^{\prime}=m-1$. Then if there exists a vertex of $(r,m^{\prime})$-rank more than $1$ in $G$, this vertex has at least $m$ neighbors, and thus is the root of a $T_{1,m}$ subgraph in $G$. For $d>1$ we proceed as follows. For a given $m$, we wish to define a suitable $m^{\prime}$. First, let $W_{d,m,r}$ denote the number of vertices of the graph obtained from $T_{d-1,m}$ by subdividing each edge $r$ times. For convenience, set $M=m\cdot W_{d,m,r}+r\cdot m+m$. Let $m^{\prime\prime}$ be the number obtained by the inductive assumption applied to d1d-1, $r$, and $M$. Then we define $m^{\prime}:=\max\{m^{\prime\prime},r\cdot m\}$. Let $v$ be a vertex of $(r,m^{\prime})$-rank more than $d$ in $G$. Then for every set $S\subseteq V(G)\setminus\{v\}$ of size at most $m^{\prime}$, the set $N_{r}^{G-S}(v)\setminus\{v\}$ contains at least one vertex of $(r,m^{\prime})$-rank at least $d$.
We claim that there exist paths $P_{1},P_{2},\ldots,P_{m}$ of length at most $r$ that are disjoint other than at $v$, and so that each path $P_{i}$ joins $v$ to a vertex $u_{i}\neq v$ that has $(r,m^{\prime})$-rank at least $d$. We find the paths greedily by adding one path at a time. Suppose that so far we have found the paths $P_{1},\ldots,P_{j}$ for some j<mj<{m}. Let $S$ be the set of all vertices $u\neq v$ that are in any of the paths $P_{1},\ldots,P_{j}$. This set $S$ contains at most $r$ vertices from each of the paths, and thus $S$ has size at most $r(m-1)\leq{m^{\prime}}$. Thus the set $N_{r}^{G-S}(v)\setminus\{v\}$ contains at least one vertex of $(r,m^{\prime})$-rank at least $d$, and we can add another path $P_{j+1}$ to our collection. This proves the claim. We note that this argument amounts to a Menger-type result about short paths, and this type of result has previously appeared in [16].
Since $m^{\prime}\geq m^{\prime\prime}$, the $(r,m^{\prime\prime})$-rank of each vertex of $G$ is at least its $(r,m^{\prime})$-rank. So by the inductive assumption applied to d1d-1, $r$, and $M$, each vertex $u_{i}$ is the root of an $\leq r$-subdivision of $T_{d-1,M}$. Let us call this $\leq r$-subdivision rooted at $u_{i}$ by $T_{i}$.
@@ 3.3 Vertex rankings and bounded expansion
Let $G$ be a graph with $\mathrm{scol}_{r}(G)=m$. Then $G$ has finite (r,m1)(r,m-1)-rank.
We stress that in Lemma 20 we are not bounding the maximum value of (r,m1)(r,m-1)-rank of any vertex of $G$, but merely claiming that the (r,m1)(r,m-1)-ranking algorithm will assign to each vertex of $G$ a finite value. This value, however, can be arbitrarily large (and in particular it is not bounded in terms of $r$ and $m$). This can be seen already on the class of trees – the class of all trees has bounded strong coloring numbers (one can set $m:=r+1$), but on the complete $m$-ary tree of depth $d$ the (r,m1)(r,m-1)-ranking algorithm will need $d$ rounds to finish.
@@
Assume for contradiction that there exists a vertex of $G$ of (r,m1)(r,m-1)-rank $\infty$. Fix an order $\leq$ on $V(G)$ that certifies that $\mathrm{scol}_{r}(G)=m$. Let $v$ be the smallest vertex of $G$ with respect to $\leq$ that has (r,m1)(r,m-1)-rank $\infty$. Let $A$ be the set of all vertices below $v$ in $\leq$, and let $i$ be the maximum (r,m1)(r,m-1)-rank of any vertex in $A$. We claim that in round $(i+1)$ of the ranking algorithm vertex $v$ received rank (meaning it has rank $i+1$), which is a contradiction with our assumption on $v$. To see this, let $S$ be the set of all vertices that are strongly $r$-reachable from $v$ in $G$. Note that every path of length at most $r$ from $v$ to a vertex $w\geq v$ must contain a vertex in $S$. It follows that $N_{r}^{G-S}(v)\setminus\{v\}$ contains only vertices of finite rank, a contradiction. ∎
@@ 4 A lemma about edge-stable graphs
Set $X_{t}:=S$. To show the existence of the remaining vertices $w_{1},\ldots,w_{t-1}$ and sets $X_{1},\ldots,X_{t-1}$, we will use the induction hypothesis with t1t-1 applied to the path $P^{\prime}:=v_{1}\ldots v_{q}$ and the set $S^{\prime}:=S\cap N^{G}(v_{q})$ (this is justified since $S^{\prime}\subseteq N^{G}(v_{q})$, the size of $S^{\prime}$ is as required, and $S^{\prime}\subseteq S$, so $S^{\prime}$ is disjoint from $N^{G}(v_{1})$). This yields vertices $w_{1},\ldots,w_{t-1}$ and sets $X_{1},\ldots,X_{t-1}$ with properties 1-3 above satisfied for $P^{\prime}$ and $S^{\prime}$. We now verify that $w_{1},\ldots,w_{t}$ together with $X_{1},\ldots,X_{t}$ satisfy properties 1-3 for $P$ and $S$. Let $i\in[t]$ be arbitrary.
@@ 5.1 Classes of tree rank at most 2
For the other direction, suppose that for every $r\in\mathbb{N}$ there exists $m\in\mathbb{N}$ such that every graph in $\mathcal{C}$ has $(r,m)$-rank at most $2$. Let $r\in\mathbb{N}$, and set $f(r):=(m+1)^{r}$ and $d(r):=m$. Let $v\in V(G)$ be arbitrary, and let $S$ be the set of vertices with $(r,m)$-rank exactly $2$ in $N_{r}^{G}(v)$ (these are the vertices of degree more than $m$). Going for a contradiction, we may assume that $|S|>f(r)$, since otherwise we are done. Thus, by Lemma 30 applied to $r$, $t:=m+1$ and $S$ in the subgraph $H$ of $G$ induced on $N_{r}^{G}(v)$, there exists a vertex $u\in N_{r}^{G}(v)$ such that there are $m+1$ vertices in $S$ that are distinct from $u$ and can be reached from $u$ by internally vertex-disjoint paths of length at most $r$. As all of the vertices in $S$ have $(r,m)$-rank 22, this shows that the $(r,m)$-rank of $u$ is more than $2$, which is a contradiction. Thus $\mathcal{C}$ is of locally almost bounded degree with functions $f$ and $d$, as desired. ∎
@@ 5.2 The case of interpretations of bounded range
Assume for contradiction that such a graph $G\in\mathcal{C}$ exists. Let $u_{1},\ldots,u_{t},w_{1},\ldots,w_{t}$ be the vertices of a half-graph of order $t$ that $G$ contains as a semi-induced subgraph. Then all of these vertices $u_{1},\ldots,u_{t},w_{1},\ldots,w_{t}$ are in $N_{2}^{G}(u_{1})$. Also, for any $i,j\in\{1,2,\ldots,t\}$ with ij>k(2)i-j>k(2), we know that $u_{i}$ and $u_{j}$ are not $k(2)$-near twins, since their adjacency differs in vertices $w_{j},w_{j+1},\ldots,w_{i-1}$. Therefore, the vertices $u_{1},u_{k(2)+2},u_{2k(2)+3},\ldots u_{m(2)\cdot k(2)+m(2)+1}$ form a set of $m(2)+1$ pairwise not $k(2)$-near-twins, a contradiction with $\mathcal{C}$ being locally almost near-covered. ∎
@@
    every vertex of $A$ is adjacent to all but at most $2k$ vertices of $B$ and every vertex of $B$ is adjacent to all but at most $2k$ vertices of AA
@@
By the definition of mutually heavy parts, the vertex $v$ has some neighbor $u^{\prime}$ that is in $A$. Since $u$ and $u^{\prime}$ each have at most $2h$ non-neighbors in $B$ and $|B|>4h$, they have a common neighbor. It follows that $\mathrm{dist}_{G}(u,v)\leq 3$.
@@
Now let us define an auxiliary graph $\widehat{G}$ where each vertex of $\widehat{G}$ is a tuple of the form $(A_{i},B_{i})$, and two distinct vertices $(A_{i},B_{i})$ and $(A_{j},B_{j})$ are adjacent if $A_{i}\cup B_{i}$ has non-empty intersection with $A_{j}\cup B_{j}$. So this graph $\widehat{G}$ has exactly $t$ vertices. Moreover, it has maximum degree less than $4(m(3)+1)$; for each vertex $(A_{i},B_{i})$, there are at most $2m(3)+1$ ways of choosing a tuple $(A,B)$ of mutually heavy parts of $\mathcal{F}$ so that either $A=A_{i}$ or $B=A_{i}$, and likewise with $B_{i}$ in place of $A_{i}$. This is using Claim Claim which implies that each vertex of $G/\mathcal{F}$ is mutually heavy with at most $m(3)$ other parts of $\mathcal{F}$ (and the extra +1+1 comes from the fact that a vertex may be mutually heavy with itself).
@@
Next, we claim that there exist $X^{\prime}\subseteq X$ and $Y^{\prime}\subseteq Y$ each of size $5h+1$ so that $X^{\prime}$ is contained in a single part $A\in\mathcal{F}$ and $Y^{\prime}$ is contained in a single part $B\in\mathcal{F}$. By symmetry, we just need to show this for $Y$. So, let $x\in X$. By Claim Claim applied to the part of $\mathcal{F}$ that contains $x$, we find that $Y$ is contained in the union of at most $m(3)$ parts of $\mathcal{F}$. So indeed, the desired set $Y^{\prime}$ exists by the pigeonhole principle.
Now, the parts $A$ and $B$ that contain $X^{\prime}$ and $Y^{\prime}$ have size at least $5h+1$, and there is a vertex in $A$ that has more than $2h$ neighbors in $B$. However, this contradicts Corollary 41 and our choice of when to flip between two parts $A$ and $B$.
@@
Finally, by Claim Claim, the set $A$ has at most $m(3)$ neighbors in the quotient graph $G/\mathcal{F}$ that are not light. Since $y$ has at most one neighbor in $\mathcal{S}(G)$ that is not a vertex of $G$, it follows that there exists $N_{y}^{\prime}\subseteq N_{y}$ of size $(|N_{y}|-k(6r)-1)/m(3)=5h+1$ so that $N_{y}^{\prime}$ is all contained in a single part $B$ of $\mathcal{F}$. Clearly this part $B$ has at least $5h+1$ vertices, and so by the construction of $\mathcal{S}(G)$ vertex $y$ from $A$ cannot have more than $2h$ neighbors in $B$, a contradiction.